OpenStudy (anonymous):

Core Focus: Absolute Value and Distance help me plz Points V, W, X, Y, and Z are graphed on a number line. The coordinate of point V is –8. What points are at a distance of 5 units from point V? Choose all answers that are correct. A. W = –13 B. X = –8 C. Y = –3 D. Z = 3

OpenStudy (welshfella):

add and subtract 5 from -8 -8 - 5 = -8 + 5 = can you work these out?

OpenStudy (welshfella):